40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 18 20:17:17 2014, Jenne, Update, LSC, ALS noisy ALS_1kHzRMSnoise.pdfALS_oolNoise.pdf
    Reply  Mon Apr 21 22:53:16 2014, manasa, Update, LSC, ALS noise ALSX_FreeInLoop.jpgALSXY_inLoop.jpg
       Reply  Tue Apr 22 01:11:42 2014, Jenne, Update, LSC, TRY 60Hz noise TRY_60Hz_noise_21Apr2014.pdf
          Reply  Tue Apr 22 21:54:50 2014, manasa, Update, LSC, TRY 60Hz noise 60HzTRY.jpg
             Reply  Tue Apr 22 22:49:10 2014, rana, Update, LSC, TRY 60Hz noise 
                Reply  Wed Apr 23 19:58:00 2014, manasa, Update, LSC, TRY 60Hz noise 
                   Reply  Tue May 6 18:32:14 2014, steve, Update, LSC, TRY 60Hz noise hunt ETMY-ISCT_EISOL.jpg
                      Reply  Wed May 7 14:01:36 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Thu May 8 17:25:17 2014, steve, Update, LSC, TRY 60Hz noise hunt 
                         Reply  Fri May 16 16:08:12 2014, steve, Update, LSC, Isolating base plates ISObaseplates.jpg
                Reply  Wed Apr 23 23:48:30 2014, manasa, Update, LSC, Y end whitening board IMG_1378.JPGIMG_1379.JPGIMG_1378.JPG
                   Reply  Thu Apr 24 14:23:09 2014, not manasa, Update, LSC, Y end whitening board 
                      Reply  Thu Apr 24 23:55:31 2014, Koji, Update, LSC, Y end whitening board P4245550.JPG
                         Reply  Fri Apr 25 10:43:57 2014, Koji, Update, LSC, (Fixed) Y end whitening board D990399-B_40m.pdfP4245552.JPGP4245553.JPGP4245551.JPG
                            Reply  Thu Dec 11 18:12:46 2014, ericq, Update, LSC, (Fixed) Y end whitening board 
                               Reply  Thu Dec 11 22:44:23 2014, rana, Update, LSC, QPD screens 
          Reply  Thu Apr 24 00:11:35 2014, Jenne, Update, LSC, Yend shutter back. 
       Reply  Sun Apr 27 20:26:19 2014, Koji, Update, LSC, Phase Tracker servo characterization ALSX_PTTF.pdfALSY_PTTF.pdf
          Reply  Sun Apr 27 21:30:59 2014, Koji, Update, LSC, ALS servo characterization ALSX_OLTF.pdfALSY_OLTF.pdf
             Reply  Mon Apr 28 10:24:10 2014, Koji, Update, LSC, error signal characterization ALSX_SPE.pdfALSY_SPE.pdf
                Reply  Mon Apr 28 10:34:51 2014, Koji, Update, LSC, new ALS servo design ALSX_OLTF_new2.pdfALSY_OLTF_new2.pdf
                   Reply  Mon Apr 28 10:48:48 2014, Koji, Update, LSC, new ALS servo design: comparison ALSX_OLTF_new.pdfALSY_OLTF_new.pdf
                      Reply  Mon Apr 28 10:59:54 2014, Koji, Update, LSC, New ALS servo design: expected error signals ALSX_SPE_new.pdfALSY_SPE_new.pdf
                         Reply  Mon Apr 28 11:08:11 2014, Koji, Update, LSC, New ALS servo design: expected error signals 140421_ALS_servo.zip
Message ID: 9838     Entry time: Tue Apr 22 01:11:42 2014     In reply to: 9836     Reply to this: 9841   9845
Author: Jenne 
Type: Update 
Category: LSC 
Subject: TRY 60Hz noise 

Quote:

 

P.S. I realigned the Y green to the arm and brought GTRY to 0.93

This evening, I was not able to successfully transition CARM from ALS to 1/sqrt(trans) signals.  The TRY time series looked odd, so I took a spectra, and we have huge 60Hz noise in TRY. 

I found a lock stretch from around 6:30pm that did not show the 60Hz noise, and then there was a lock stretch around 8pm that did have the noise.  So, something happened at the Yend between 6:30 and 8pm tonight.

Asking around, this was the time frame in which Manasa was down at the Yend to realign the green beam, and to check cabling for the PZT_OUT and ERR_MON signals to the ADC.

Looking at the spectra, Rana noted that we have even as well as odd harmonics of the 60Hz line, which is unusual.

TRY_60Hz_noise_21Apr2014.pdf

To try to diagnose the problem, Rana and I tried to make sure no cables' connectors were touching, and that no equipment was plugged in that shouldn't be.  We noticed that none of: the shutter, the Thorlabs TRY PD, or the QPD TRY are isolated from the table.  To see if perhaps the shutter was the problem, I turned off the power to the Yend green shutter, and unplugged the cable.  The cable is laying on the table, with the connector sitting on a piece of plastic to isolate it.  Removing the shutter from the system did not change anything. 

We don't see the 60Hz noise in the Xarm, so it's not on the laser light itself.  Also, we don't see the 60Hz lines in the Yarm feedback signal, so we're not putting the lines onto the mirror, and thus onto the Yarm's light. 

Manasa, can you please take a look, and see if you can figure out what is going on?  We need TRY so that we can transition to 1/sqrt(trans) signals for CARM.  Thanks!!

ELOG V3.1.3-